# Measuring Meaning Composition in the Human Brain with Composition Scores from Large Language Models

### v1

@@ Abstract
Changjiang Gao1 Jixing Li2 Jiajun Chen1 Shujian Huang1thanks: Corresponding author 1National Key Laboratory for Novel Software Technology, Nanjing University 2Department of Linguistics and Translation, City University of Hong Kong gaocj@smail.nju.edu.cn  jixingli@cityu.edu.hk {chenjj, huangsj}@nju.edu.cn
@@ 3.1.1 The key-value memory interpretation
where $\mathbf{x}\in\mathbb{R}^{d}$ is the input vector, $K,V\in\mathbb{R}^{d_{m}}$ are the two linear layers inside the FFN block, and $f$ is the activation function. This formulation can be viewed as a generalized expression of a neural memory Sukhbaatar et al. (2015):
@@ 3.1.3 Activation-based approximation
Figure 2: (a) Density plot of word frequency, node counts based on the top-down, bottom-up and left-corner node counts. (b) Correlation matrix among the 4 control variables.
@@ 5.2.2 Significant brain clusters
Figure 7: Significant brain clusters for Composition Scores and the significant layers from LLaMA2-chat. The light pink regions in the brain indicate the language mask. The orange and red lines depict the normalized $\beta$ value for each layer of the two models. The grey lines depict the normalized $\beta$ value for each layer of the random models. The shaded region indicates the significant layers. *** indicates $p$ <0.001.
@@ 6 Conclusion
In this paper, we introduce a novel model-based metric, the Composition Score, designed to quantify sentence-level meaning composition, and examine its correlation with human neural activity.
We identify several brain clusters significantly correlated with the Composition Score, including those associated with word frequency, syntactic structure, and general sensitivity to words. This suggests a multifaceted nature of meaning composition during human sentence comprehension.
@@ Appendix A Computational Resource
Appendix B Results of LLaMA2-Base
Figure 8 in Appendix B displays Pearson’s $r$ among the layerwise Composition Score from LLaMA2-base. Similar to LLaMA2-chat, the layers form small correlated clusters and do not exhibit high overall correlation. Figure 10 illustrates the regression scores between the layerwise Composition Score from LLaMA2-base and the control variables. The results mirror those of LLaMA2-chat. Figure 9 in Appendix B depicts the significant brain clusters correlated with the layerwise Composition Scores from LLaMA2-base. Similar to LLaMA2-chat, there are two separated layer clusters in the first and second half of the model layers respectively, and the brain clusters closely resemble those of LLaMA2-chat.

### v3

@@ Abstract
Changjiang Gao1 Jixing Li2thanks: Corresponding authors, equal contribution. Jiajun Chen1 Shujian Huang1footnotemark: 1National Key Laboratory for Novel Software Technology, Nanjing University 2Department of Linguistics and Translation, City University of Hong Kong gaocj@smail.nju.edu.cn  jixingli@cityu.edu.hk {chenjj, huangsj}@nju.edu.cn
@@ 3.1.1 The key-value memory interpretation
where $\mathbf{x}\in\mathbb{R}^{d}$ is the input vector, $K,V\in\mathbb{R}^{d_{m}\times d}$ are the two linear layers inside the FFN block, and $f$ is the activation function. This formulation can be viewed as a generalized expression of a neural memory Sukhbaatar et al. (2015):
@@ 3.1.3 Activation-based approximation
Figure 3: (a) Density plot of word frequency, node counts based on the top-down, bottom-up and left-corner node counts. Note that density plot is different from a histogram such that values on the y-axis here represent probability density and the total area under the curve integrates to one. (b) Correlation matrix among the 4 control variables.
@@ 5.2.2 Significant brain clusters
Figure 7: Significant brain clusters for Composition Scores and the significant layers from LLaMA2-chat. The light pink regions in the brain indicate the language mask. The grey lines depict the normalized $\beta$ value for each layer of the random models. The shaded region indicates the significant layers. *** indicates $p$ <0.001.
@@ 5.3.2 Significant brain clusters
5.4 Comparison with fMRI results for hidden layer activity
The results presented above identify significant brain clusters associated with our proposed Composition Score. However, it remains to be determined whether this score provides additional information beyond that obtained through typical encoding models that utilize hidden layer activity Huth et al. (2016); Jain and Huth (2018); Kay et al. (2008); Naselaris et al. (2011). To explore this, we conduct further analyses using hidden layer activity from LLaMA2-chat and base models as features in our encoding model. We first compress the high-dimensional hidden states into 100-dimensional vectors using PCA to save computational resources. Subsequently, we perform ridge regression to predict the activity of each vertex within a language mask for individual subjects. We then identify significant brain clusters and layers for the normalized regression scores ($R^{2}$) at the group level using the same spatiotemporal clustering analysis outlined in Section 3.3.
The analysis reveals a significant cluster within a broad left temporal-parietal network (N vertices = 515, t = 1.68, p < 0.0001) for the chat model (see Figure 8). Similar findings are observed for the base model (detailed in Appendix C). The clustering patterns differ markedly from those associated with the Composition Score, suggesting that the model’s hidden states and our Composition Score capture different aspects of meaning comprehension.
Figure 8: Significant brain clusters for hidden layer activity and the significant layers from LLaMA2-chat. The light pink regions in the brain indicate the language mask. The grey line depicts the normalized $R^{2}$ value for the random model.
6 Discussion
6.1 Composition Score vs. hidden states of LLMs
The Composition Score proposed in this paper adopts the quotient of the minimum and maximum distances between the vocabulary distributions between the individual neural memory and the layer output. Our metric emphasizes the "degree of compositionality," specifically assessing whether the final model output is dependent on a limited subset of neurons or a broader, distributed network. This approach is distinct from previous studies that compare the hidden layer activity of language models with brain activity during naturalistic comprehension Huth et al. (2016); Goldstein et al. (2022); Schrimpf et al. (2021); Caucheteux and King (2022). As shown in Figure 8, the hidden states of LLaMA2-chat correspond to a large cluster along the left temporal-parietal pathway, contrasting with the activities in the LIFG and LaSTG observed for the Composition Score (see Figure 7). This suggests that while the hidden layer activity may represent the unfolding of meaning in general, our Composition Score provides insights into the process of meaning composition.
6.2 Composition Score in LLMs vs. meaning composition in the brain
The term "memory composition," was used in Geva et al. (2021) to explain the function of FFN blocks in Transformers, and we consider it to potentially mirror the operations for meaning composition in the human brain. Previous neuroimaging studies have identified the LATL and LPTL as crucial for meaning composition Bemis and Pylkkänen (2011, 2013); Parrish and Pylkkänen (2022); Zhang and Pylkkänen (2015); Flick and Pylkkänen (2020); Hagoort (2005); Lyu et al. (2019); Matchin et al. (2019); Matchin and Hickok (2020); Li and Pylkkänen (2021), but a mechanistic understanding of how complex meanings are assembled from individual neurons remains elusive. Our Composition Score quantifies the degree of composition in LLMs for each word within a context, with a high score indicating the involvement of more neurons in the representation of a word. This concept also aligns with literature on "sparse" or "distributed" coding for object representation Quiroga et al. (2008) such as the "Jennifer Aniston neuron" study by Quiroga et al. (2005), which showed that individual neurons in the medial temporal lobe (MTL) can be selectively activated by different images of the same individual.
@@ 7 Conclusion
In this paper, we introduce a novel model-based metric, the Composition Score to quantify meaning composition. We examine its correlation with human neural activity and identify several brain clusters, offering insights into the process of meaning composition in the brain.
@@
Moreover, The Composition Score proposed in this paper adopts the quotient of the minimum and maximum distances between the vocabulary distributions between the individual neural memory and the layer output, which tries to capture the populational variance of the memorized and predicted vocabulary distributions. However, this metric may over-simplify the true populational variance, especially when the distances are rather sharply than evenly distributed. In such cases, the calculation of the Composition Scores can be adapted to include more detailed features of the distance distribution, by using information theory-based metrics such as entropy or the Gini Index.
@@
Acknowledgements
We would like to thank the anonymous reviewers for their insightful comments. Shuanjian Huang and Jixing Li are the corresponding authors who contribute equally to this work. This work is supported by National Science Foundation of China (No. 62376116, 62176120), the Liaoning Provincial Research Foundation for Basic Research (No. 2022-KF-26-02), research project of Nanjing University-China Mobile Joint Institute.
@@ Appendix A Computational Resource
Appendix B Results of LLaMA2-base for Composition Score
Figure 10 in Appendix B displays Pearson’s $r$ among the layerwise Composition Score from LLaMA2-base. Similar to LLaMA2-chat, the layers form small correlated clusters and do not exhibit high overall correlation. Figure 9 illustrates the regression scores between the layerwise Composition Score from LLaMA2-base and the control variables. The results mirror those of LLaMA2-chat. Figure 11 in Appendix B depicts the significant brain clusters correlated with Composition Scores from LLaMA2-base. Similar to LLaMA2-chat, there are two separated clusters in the first and second half of the model layers respectively, and the brain clusters closely resemble those of LLaMA2-chat.
@@ Appendix B Results of LLaMA2-base for Composition Score
Figure 12: Significant brain clusters for hidden layer activity and the significant layers from LLaMA2-base. The light pink regions in the brain indicate the language mask. The grey line depicts the normalized $R^{2}$ value for the random model.
Appendix C Results of LLaMA2-base for hidden layer activity
Figure 12 in Appendix C depicts the significant brain clusters correlated with hidden layer activity from LLaMA2-base. Similar to LLaMA2-chat, there is a wide left temporal-parietal network significantly associated with the hidden states for all layers (N vertices = 503, t = 1.69, p < 0.0001).